# Sequential Monte-Carlo testing by betting

### v4

@@ 6 Every permutation test can be recovered by betting
$E_{t-1}(1)+\ldots+E_{t-1}(t)$ $=t,$ (12)
where $B_{t|r}[L]$ (or $B_{t|r}[W]$) denotes our bet on a loss (or win) at step $t$ if $\mathrm{rank}(Y_{0};Y_{0}^{t-1})=r$. It makes sense that we focus on such betting strategies as the e-value $E_{t}$ solely depends on $\mathrm{rank}(Y_{0};Y_{0}^{t})$. The equations (9) and (10) ensure that $E_{t-1}\cdot B_{t}$ leads to the same e-value as $E_{t}$ and the equations (11) and (12) ensure that the sum of the betting vectors $B_{t}$ and $E_{t-1}$ is $t+1$ and $t$, respectively.
@@
The unique solution of the system of equations (9)-(12) is given by
$E_{t-1}(r)$ $=\frac{r}{t+1}E_{t}(r+1)+\frac{t+1-r}{t+1}E_{t}(r)$
$B_{t|r}[W]$ $=\frac{E_{t}(r)}{E_{t-1}(r)}$
$B_{t|r}[L]$ $=\frac{E_{t}(r+1)}{E_{t-1}(r)}$
for all $r=1,\ldots,t$ with the convention $/0=0$ 0. Furthermore, every e-value $E_{t}(\mathrm{rank}(Y_{0};Y_{0}^{t}))$ can be obtained by a sequential betting strategy $E_{t}(\mathrm{rank}(Y_{0};Y_{0}^{t}))=\prod_{s=1}^{t}B_{s}(R_{s})$, where
$B_{s}(r)=\begin{cases}B_{s|\mathrm{rank}(Y_{0};Y_{0}^{s-1})}[L],\text{ if }r\leq\mathrm{rank}(Y_{0};Y_{0}^{s-1})\\ B_{s|\mathrm{rank}(Y_{0};Y_{0}^{s-1})}[W],\text{ otherwise }\end{cases}$
with $B_{s|\mathrm{rank}(Y_{0};Y_{0}^{s-1})}[L]$ and $B_{s|\mathrm{rank}(Y_{0};Y_{0}^{s-1})}[W]$ being defined recursively by the formulas above.
@@
We solve for equations (9)-(11) and then show that (12) is automatically fulfilled as well. Equations (10) and (11) immediately give $B_{t}[W]=\frac{E_{t}(r)}{E_{t-1}(r)}$ and $B_{t|r}[L]=\frac{t+1-(t+1-r)B_{t|r}[W]}{r}$. Inserting this into (9) yields
$E_{t-1}(r)\left(\frac{t+1-(t+1-r)\frac{E_{t}(r)}{E_{t-1}(r)}}{r}\right)=E_{t}(r+1)$
@@
$\sum_{r=1}^{t}E_{t-1}(r)$ $=\sum_{r=1}^{t}\frac{r}{t+1}E_{t}(r+1)+\sum_{r=1}^{t}\frac{t+1-r}{t+1}E_{t}(r)$
$=\frac{t}{t+1}E_{t}(t+1)+\frac{t}{t+1}E_{t}(1)+\sum_{r=2}^{t}\frac{r-1}{t+1}E_{t}(r)+\sum_{r=2}^{t}\frac{t+1-r}{t+1}E_{t}(r)$
$=\frac{t}{t+1}E_{t}(t+1)+\frac{t}{t+1}E_{t}(1)+\frac{t}{t+1}\sum_{r=2}^{t}E_{t}(r)$
$=\frac{t}{t+1}\sum_{r=1}^{t+1}E_{t}(r)=t,$
@@
This theorem shows that every e-value that is a function of $\mathrm{rank}(Y_{0};Y_{0}^{t})$ for a fixed $t$ can also be obtained by our sequential betting strategy. Therefore, also every level-$\alpha$ test $\phi_{\alpha}(\mathrm{rank}(Y_{0};Y_{0}^{t}))$, where $\phi_{\alpha}(\mathrm{rank}(Y_{0};Y_{0}^{t}))$ takes values in $\{0,1\}$ and $\mathbb{P}_{H_{0}}(\phi_{\alpha}(\mathrm{rank}(Y_{0};Y_{0}^{t}))=1)\leq\alpha$, can be obtained (or uniformly improved) by our sequential betting strategy, since $E_{t}^{\alpha}=\phi_{\alpha}(\mathrm{rank}(Y_{0};Y_{0}^{t}))/\alpha$ is a valid e-value. This is captured in the following corollary.
@@
For every level-$\alpha$ test $\phi_{\alpha}(\mathrm{rank}(Y_{0};Y_{0}^{t}))$ that only depends on $\mathrm{rank}(Y_{0};Y_{0}^{t})$ for some $t\geq 1$, there is a betting strategy $B_{1},\ldots,B_{t}$ such that $1\{W_{t}\geq 1/\alpha\}\geq\phi_{\alpha}(\mathrm{rank}(Y_{0};Y_{0}^{t}))$ almost surely, where $W_{t}=\prod_{s=1}^{t}B_{s}(R_{s})$.
@@
Note that every p-value $\mathfrak{p}$ gives rise to a family of level-$\alpha$ tests $\boldsymbol{\phi}^{\mathfrak{p}}=(\phi_{\alpha}^{\mathfrak{p}})_{\alpha\in[0,1]}$ by $\phi_{\alpha}^{\mathfrak{p}}=1\{\mathfrak{p}\leq\alpha\}$. Thus, if each of these $\phi_{\alpha}^{\mathfrak{p}}$ depends only on $\mathrm{rank}(Y_{0};Y_{0}^{t})$ for some $t\geq 1$, we could derive sequential versions of these level-$\alpha$ tests and then calibrate them into an anytime-valid p-value.
@@
For example, let $B\in\mathbb{N}$ and $\alpha\in[0,1]$ be fixed and consider the e-value $E_{B}^{\alpha}(\mathrm{rank}(Y_{0};Y_{0}^{B-1}))$ with the betting vector $E_{B}^{\alpha}$ given by
$1/\alpha,\ldots,1/\alpha,a,0,\ldots,0,$ (13)
@@
where the switch from $1/\alpha$ to $a$ happens after index $\lfloor(B+1)\alpha\rfloor$ and $a$ is a constant $0\leq a<1/\alpha$ such that the sum of $E_{B}^{\alpha}$ equals $B$. If $(B+1)\alpha$ is a whole number, $a$ equals 00. Theorem 6.1 shows that this e-value can be obtained by our sequential betting strategy $E_{B}^{\alpha}(\mathrm{rank}(Y_{0};Y_{0}^{B}))=\prod_{t=1}^{B}B_{t}^{\alpha,B}(R_{t})$ for some betting vectors $B_{1}^{\alpha,B},\ldots,B_{B}^{\alpha,B}$. Note that $E_{B}^{\alpha}(\mathrm{rank}(Y_{0};Y_{0}^{B}))$ yields the same decisions as the permutation p-value $\mathfrak{p}^{\mathrm{perm}}_{B}$ at level $\alpha$, since $\mathfrak{p}^{\mathrm{perm}}_{B}\leq\alpha$ iff $\mathrm{rank}(Y_{0};Y_{0}^{B})\leq(B+1)\alpha$. Therefore,
@@
$\mathfrak{p}_{t}^{\mathrm{av},B}=\inf\left\{\alpha\in[0,1]\Bigg{|}\exists r=1,\ldots,t:\prod_{s=1}^{r}B_{s}^{\alpha,B}(R_{s})\geq 1/\alpha\right\}\quad t=1,\ldots,B$ (14)
@@
is an anytime-valid generalization of the permutation p-value, since $\mathfrak{p}_{B}^{\mathrm{av},B}\leq\alpha$ iff $\mathrm{rank}(Y_{0};Y_{0}^{B})\leq(B+1)\alpha$ but $\mathfrak{p}_{t}^{\mathrm{av},B}$ is valid for every stopping time $\tau\leq B$. Indeed, we could even continue sampling after step $B$. However, if $a=0$, the decisions made by our betting strategies would not change (as we either hit $1/\alpha$ or 00 at step $B$) and thus $\mathfrak{p}_{t}^{\mathrm{av},B}$ would not either. To see how this p-value looks like if we stop at some time $\tau\leq B$, note that $\prod_{s=1}^{\tau}B_{s}^{\alpha,B}(R_{s})=E_{\tau}^{\alpha,B}(\mathrm{rank}(Y_{0};Y_{0}^{\tau}))$, where $E_{\tau}^{\alpha,B}$ can be calculated recursively by $E_{t-1}^{\alpha,B}(r)=\frac{r}{t+1}E_{t}^{\alpha,B}(r+1)+\frac{t+1-r}{t+1}E_{t}^{\alpha,B}(r)$ with the starting point $E_{B}^{\alpha,B}=E_{B}^{\alpha}$ (see Theorem 6.1). Since $E_{B}^{\alpha}(r)=1/\alpha$ for all $r\leq\lfloor(B+1)\alpha\rfloor$ and $E_{B}^{\alpha}(r)<1/\alpha$ otherwise, this shows that $E_{\tau}^{\alpha,B}$ is given by
$1/\alpha,\ldots,1/\alpha,a_{1},\ldots,a_{B-\tau+1},0,\ldots,0,$
@@
where $a_{1},\ldots,a_{B-\tau+1}$ are constants $0\leq a_{i}<1/\alpha$ and the switch from $1/\alpha$ to $a_{1}$ happens after index $\lfloor(B+1)\alpha\rfloor-(B-\tau)$. Hence, $E_{\tau}^{\alpha}\geq 1/\alpha$ iff $\mathrm{rank}(Y_{0},Y_{0}^{\tau})\leq(B+1)\alpha-(B-\tau)$, which shows that $\mathfrak{p}_{\tau}^{\mathrm{av},B}=(\mathrm{rank}(Y_{0},Y_{0}^{\tau})+B-\tau)/(B+1)$ for $\tau\leq B$. This is the permutation p-value $\mathfrak{p}^{\mathrm{perm}}_{B}$ we would obtain if all permutations after step $\tau$ were losses.
@@
In the same way one could recover the Besag-Clifford p-value $\mathfrak{p}^{\mathrm{BC}}$. For this, consider again the betting vector $E_{B}^{\alpha}$ defined in (13) but let $B=\min\left(B_{\max},\lceil h/\alpha\rceil-1\right)$, where $h$ and $B_{\max}$ are the predefined number of losses and number of permutations for the Besag-Clifford method, and define
@@
$\mathfrak{p}_{t}^{\mathrm{av},B_{\max},h}=\inf\left\{\alpha\in[0,1]\Bigg{|}\exists r=1,\ldots,t:\prod_{s=1}^{\min\left(r,B_{\max},\lceil h/\alpha\rceil-1\right)}B_{s}^{\alpha,\min\left(B_{\max},\lceil h/\alpha\rceil-1\right)}(R_{s})\geq 1/\alpha\right\}.$ (15)
@@
Let $\tau\leq\gamma(h,B_{\max})$ be any stopping time. Since $\mathrm{rank}(Y_{0};Y_{0}^{\tau-1})\leq h$, we have $E_{\lceil h/\alpha\rceil-1}^{\alpha,\min\left(B_{\max},\lceil h/\alpha\rceil-1\right)}(\mathrm{rank}(Y_{0};Y_{0}^{\lceil h/\alpha\rceil-1}))=1/\alpha$ for all $\alpha\in[0,1]$ with $\tau>\lceil h/\alpha\rceil-1$, meaning we reject at level $\alpha$ if sampling $\lceil h/\alpha\rceil-1$ permutations before reaching $h$ losses. Now consider $\alpha\in[0,1]$ such that $\tau\leq\min(B_{\max},\lceil h/\alpha\rceil-1)$. With the same reasoning as above, it follows that $E_{\tau}^{\alpha,\min(B_{\max},\lceil h/\alpha\rceil-1)}\geq 1/\alpha$ iff
$\mathrm{rank}(Y_{0};Y_{0}^{\tau})\leq(\min(B_{\max},\lceil h/\alpha\rceil-1)+1)\alpha-(\min(B_{\max},\lceil h/\alpha\rceil-1)-\tau).$
@@
If $B_{\max}<\lceil h/\alpha\rceil-1$, this is equivalent to $(\mathrm{rank}(Y_{0},Y_{0}^{\tau})+(B-\tau))/(B+1)\leq\alpha$ and if $B_{\max}\geq\lceil h/\alpha\rceil-1$, it is equivalent to $h/(\tau+1+h-\mathrm{rank}(Y_{0};Y_{0}^{\tau}))\leq\alpha$. Altogether, this implies that
$\mathfrak{p}_{\tau}^{\mathrm{av},B_{\max},h}=\min\left(\frac{h}{\tau+1+h-\mathrm{rank}(Y_{0};Y_{0}^{\tau})},\frac{\mathrm{rank}(Y_{0};Y_{0}^{\tau})+(B_{\max}-\tau)}{B_{\max}+1}\right)\quad(\tau\leq\gamma(h,B_{\max}))$
@@
and therefore $\mathfrak{p}_{t}^{\mathrm{av},B_{\max},h}$ is an anytime-valid generalization of the p-value by Besag and Clifford [3]. Again, $\mathfrak{p}_{\tau}^{\mathrm{av},B_{\max},h}$ is the same as the Besag-Clifford $p$-value $\mathfrak{p}^{\mathrm{BC}}_{\gamma(h,B_{\max})}$ if we would only observe losses after step $\tau$; and if we would continue sampling after $\gamma(h,B)$, the p-value $\mathfrak{p}_{\tau}^{\mathrm{av},B_{\max},h}$ would not change. This generalizes our results from Section 2.1, where we showed that the aggressive strategy recovers the Besag-Clifford method for $h=1$, to general $h$. Indeed, $\mathfrak{p}_{t}^{\mathrm{av},\infty,1}=\mathfrak{p}^{\mathrm{agg}}_{t}$ for all $t\in\mathbb{N}$.
@@
It is interesting how the permutation p-value $\mathfrak{p}^{\mathrm{perm}}$ and the Besag-Clifford p-value $\mathfrak{p}^{\mathrm{BC}}$ arise from our approach. Indeed, the bet $B^{B}_{\alpha}$ on $\mathrm{rank}(Y_{0};Y_{0}^{t})$ in (13) is the most powerful bet we could make for fixed $\alpha$ and number of permutations $B$. Therefore, the resulting sequential betting strategy $B_{1}^{\alpha,B},\ldots,B_{B}^{\alpha,B}$ is optimal if $\alpha$ and $B$ are fixed. However, we claim that fixing the number of permutations is not the best way to perform permutation testing, since it it is more reasonable to draw a larger number of permutations if the decision is tight and fewer permutations if the decision is unambiguous. This is provided by our log-optimality based strategies.

### v5

@@ 6.1 Recursive betting strategy for the construction of any permutation e-value
$E_{t-1}(0)+\ldots+E_{t-1}(t-1)$ $=t.$ (16)
The equations (13) and (14) ensure that $E_{t-1}\cdot B_{t}$ leads to the same e-value as $E_{t}$ and the equations (15) and (16) ensure that $B_{t}$ and $E_{t-1}$ are valid betting functions according to (2) and (12), respectively.
@@
Every e-value $E_{t}(L_{t})$ can be obtained by a sequential betting strategy $E_{t}(L_{t})=\prod_{r=1}^{t}B_{r|L_{r-1}}(I_{r})$, where $B_{r|\ell}(0)$ and $B_{r|\ell}(1)$ are defined recursively via
$E_{t-1}(\ell)$ $=\frac{\ell+1}{t+1}E_{t}(\ell+1)+\frac{t-\ell}{t+1}E_{t}(\ell)$
$B_{t|\ell}(0)$ $=\frac{E_{t}(\ell)}{E_{t-1}(\ell)}$
$B_{t|\ell}(1)$ $=\frac{E_{t}(\ell+1)}{E_{t-1}(\ell)},$
which are the unique solutions to the equations (13)-(16). We use the convention $/0=0$ 0.
@@
We solve for equations (13)-(15) and then show that (16) is automatically fulfilled as well. Equations (14) and (15) immediately give $B_{t|\ell}(0)=\frac{E_{t}(\ell)}{E_{t-1}(\ell)}$ and $B_{t|\ell}(1)=\frac{t+1-(t-\ell)B_{t|\ell}(0)}{\ell+1}$. Inserting this into (13) yields
$E_{t-1}(\ell)\left(\frac{t+1-(t-\ell)\frac{E_{t}(\ell)}{E_{t-1}(\ell)}}{\ell+1}\right)=E_{t}(\ell+1)$
@@
$\sum_{\ell=0}^{t-1}E_{t-1}(\ell)$ $=\sum_{\ell=0}^{t-1}\frac{\ell+1}{t+1}E_{t}(\ell+1)+\sum_{\ell=0}^{t-1}\frac{t-\ell}{t+1}E_{t}(\ell)$
$=\frac{t}{t+1}E_{t}(t)+\frac{t}{t+1}E_{t}(0)+\sum_{\ell=1}^{t-1}\frac{\ell}{t+1}E_{t}(\ell)+\sum_{\ell=1}^{t-1}\frac{t-\ell}{t+1}E_{t}(\ell)$
$=\frac{t}{t+1}E_{t}(t)+\frac{t}{t+1}E_{t}(0)+\frac{t}{t+1}\sum_{\ell=1}^{t-1}E_{t}(\ell)$
$=\frac{t}{t+1}\sum_{\ell=0}^{t}E_{t}(\ell)=t,$
@@
This theorem shows that every e-value that is a function of $L_{t}$ for a fixed $t$ can also be obtained by our sequential betting strategy. Therefore, also every level-$\alpha$ test $\phi_{\alpha}(L_{t})$, where $\phi_{\alpha}(L_{t})$ takes values in $\{0,1\}$ and $\mathbb{P}_{H_{0}}(\phi_{\alpha}(L_{t})=1)\leq\alpha$, can be obtained (or uniformly improved) by our sequential betting strategy, since $E_{t}^{\alpha}(L_{t})=\phi_{\alpha}(L_{t})/\alpha$ is a valid e-value. This is captured in the following corollary.
@@
For every level-$\alpha$ test $\phi_{\alpha}(L_{t})$ that only depends on the number of losses $L_{t}$ for some $t\geq 1$, there is a betting strategy $B_{1},\ldots,B_{t}$ such that $\mathbbm{1}\{W_{t}\geq 1/\alpha\}\geq\phi_{\alpha}(L_{t})$ almost surely, where $W_{t}=\prod_{r=1}^{t}B_{r}(I_{s})$.
@@
Note that every p-value $Q$ gives rise to a family of level-$\alpha$ tests $\boldsymbol{\phi}^{Q}=(\phi_{\alpha}^{Q})_{\alpha\in[0,1]}$ by $\phi_{\alpha}^{Q}=\mathbbm{1}\{Q\leq\alpha\}$. Thus, if each of these $\phi_{\alpha}^{Q}$ depends only on $L_{t}$ for some $t\geq 1$, we could derive sequential versions of these level-$\alpha$ tests and then calibrate them into an anytime-valid p-value. In the following subsections, we use this to obtain anytime-valid generalizations of the permutation p-value (1) and the Besag-Clifford p-value (8).
6.2 Generalizing the permutation p-value by betting
@@
Let $T\in\mathbb{N}$ and $\alpha\in[0,1]$ be fixed and consider the e-value $E_{T}^{\alpha}(L_{T})$ with the betting function $E_{T}^{\alpha}$ given by
$E_{T}^{\alpha}(\ell)=\begin{cases}1/\alpha,&\ell\leq\lfloor(T+1)\alpha\rfloor-1\\ a,&\ell=\lfloor(T+1)\alpha\rfloor\\ 0,&\text{otherwise}\end{cases}\quad(\ell\in\{0,1,\ldots,T\}),$ (17)
@@
where $a$ is a constant $0\leq a<1/\alpha$ such that the sum of $E_{T}^{\alpha}(\ell)$ equals $T+1$. If $(T+1)\alpha$ is a whole number, $a$ equals 00. Theorem 6.1 shows that this e-value can be obtained by our sequential betting strategy $E_{T}^{\alpha}(L_{T})=\prod_{t=1}^{T}B_{t}^{\alpha,T}(I_{t})$ for some betting functions $B_{1}^{\alpha,T},\ldots,B_{T}^{\alpha,T}$. Define
@@
$Q_{t}^{\mathrm{av},T}:=\inf\left\{\alpha\in[0,1]\Bigg{|}\exists r=1,\ldots,t:\prod_{s=1}^{\min(r,T)}B_{s}^{\alpha,T}(I_{s})\geq 1/\alpha\right\}.$ (18)
In the following proposition, we show that $(Q_{t}^{\mathrm{av},T})_{t\in\mathbb{N}}$ is an anytime-valid generalization of the permutation p-value (1).
Proposition 6.3.
$(Q_{t}^{\mathrm{av},T})_{t\in\mathbb{N}}$ is an anytime-valid p-value with
$Q_{\tau}^{\mathrm{av},T}=(L_{\tau}+1+T-\tau)/(T+1)$
for all stopping times $\tau\leq T$ and $Q_{\tau}^{\mathrm{av},T}=Q_{T}^{\mathrm{av},T}$ otherwise. In particular, $Q_{T}^{\mathrm{av},T}=Q^{\mathrm{perm}}_{T}$, implying that $(Q_{t}^{\mathrm{av},T})_{t\in\mathbb{N}}$ is an anytime-valid generalization of the permutation p-value (1).
Proof.
@@
Ville’s inequality implies that $(Q_{t}^{\mathrm{av},T})_{t\in\mathbb{N}}$ is an anytime-valid p-value. Furthermore, note that for all $\tau\leq T$ we have $\prod_{s=1}^{\tau}B_{s}^{\alpha,T}(I_{s})=E_{\tau}^{\alpha,T}(L_{\tau})$, where $E_{\tau}^{\alpha,T}$ can be calculated recursively by $E_{t-1}^{\alpha,T}(\ell)=\frac{\ell+1}{t+1}E_{t}^{\alpha,T}(\ell+1)+\frac{t-\ell}{t+1}E_{t}^{\alpha,T}(\ell)$ with the starting point $E_{T}^{\alpha,T}=E_{T}^{\alpha}$ (see Theorem 6.1). Since $E_{T}^{\alpha}(\ell)=1/\alpha$ for all $\ell\leq\lfloor(T+1)\alpha\rfloor-1$ and $E_{T}^{\alpha}(\ell)<1/\alpha$ otherwise, this shows that $E_{\tau}^{\alpha,T}$ is given by
$E_{\tau}^{\alpha,T}(\ell)=\begin{cases}1/\alpha,&\ell\leq\lfloor(T+1)\alpha\rfloor-1-(T-\tau)\\ a_{\ell},&\lfloor(T+1)\alpha\rfloor-(T-\tau)\leq\ell\leq\lfloor(T+1)\alpha\rfloor\\ 0,&\text{otherwise}\end{cases}\quad(\ell\in\{0,1,\ldots,T\}),$
where the $a_{\ell}$ are constants $0\leq a_{\ell}<1/\alpha$. Hence, $E_{\tau}^{\alpha,T}(L_{\tau})\geq 1/\alpha$ iff $L_{\tau}+1\leq(T+1)\alpha-(T-\tau)$, which shows that $Q_{\tau}^{\mathrm{av},T}=(L_{\tau}+1+T-\tau)/(T+1)$ for $\tau\leq T$. ∎
@@
Note that $Q_{\tau}^{\mathrm{av},T}$ is the permutation p-value $Q^{\mathrm{perm}}_{T}$ we would obtain if all permutations after step $\tau$ were losses.
6.3 Generalizing the Besag-Clifford p-value by betting
@@
The Besag-Clifford p-value $Q^{\mathrm{BC}}$ (8) can be generalized in the same way as the permutation p-value. For this, consider again the betting function $E_{T}^{\alpha}$ defined in (17) but let $T=\min\left(T_{\max},\lceil h/\alpha\rceil-1\right)$, where $h$ and $T_{\max}$ are the predefined number of losses and number of permutations for the Besag-Clifford method, and define
@@
$Q_{t}^{\mathrm{av},T_{\max},h}:=\inf\left\{\alpha\in[0,1]\Bigg{|}\exists r=1,\ldots,t:\prod_{s=1}^{\min\left(r,T_{\max},\lceil h/\alpha\rceil-1\right)}B_{s}^{\alpha,\min\left(T_{\max},\lceil h/\alpha\rceil-1\right)}(I_{s})\geq 1/\alpha\right\}.$ (19)
Proposition 6.4.
$(Q_{t}^{\mathrm{av},T_{\max},h})_{t\in\mathbb{N}}$ is an anytime-valid p-value with
$Q_{\tau}^{\mathrm{av},T_{\max},h}=\min\left(\frac{h}{\tau+h-L_{\tau}},\frac{L_{\tau}+1+T_{\max}-\tau}{T_{\max}+1}\right)$
for all stopping times $\tau\leq\gamma(h,T_{\max})$ and $Q_{\tau}^{\mathrm{av},T_{\max},h}=Q_{\gamma(h,T_{\max})}^{\mathrm{av},T_{\max},h}$ otherwise. In particular, $Q_{\gamma(h,T_{\max})}^{\mathrm{av},T_{\max},h}=Q^{\mathrm{BC}}_{\gamma(h,T_{\max})}$, implying that $(Q_{t}^{\mathrm{av},T})_{t\in\mathbb{N}}$ is an anytime-valid generalization of the Besag-Clifford p-value (8).
Proof.
@@
Ville’s inequality implies that $(Q_{t}^{\mathrm{av},T_{\max},h})_{t\in\mathbb{N}}$ is an anytime-valid p-value. Furthermore, let $\tau\leq\gamma(h,T_{\max})$ be any stopping time. First, consider $\alpha\in[0,1]$ such that $\lceil h/\alpha\rceil-1<\tau$, implying that $L_{\lceil h/\alpha\rceil-1}\leq h-1$ and $T_{\max}>\lceil h/\alpha\rceil-1$. Then, we have $E_{\min\left(T_{\max},\lceil h/\alpha\rceil-1\right)}^{\alpha}(L_{\lceil h/\alpha\rceil-1})=1/\alpha$. Now consider $\alpha\in[0,1]$ such that $\tau\leq\min(T_{\max},\lceil h/\alpha\rceil-1)$. With the same reasoning as in the proof of Proposition 6.3, it follows that $E_{\tau}^{\alpha,\min(T_{\max},\lceil h/\alpha\rceil-1)}(L_{\tau})\geq 1/\alpha$ iff
$L_{\tau}+1\leq(\min(T_{\max},\lceil h/\alpha\rceil-1)+1)\alpha-(\min(T_{\max},\lceil h/\alpha\rceil-1)-\tau).$
@@
If $T_{\max}<\lceil h/\alpha\rceil-1$, this is equivalent to $(L_{\tau}+1+T_{\max}-\tau)/(T_{\max}+1)\leq\alpha$ and if $T_{\max}\geq\lceil h/\alpha\rceil-1$, it is equivalent to $h/(\tau+h-L_{\tau}))\leq\alpha$. Altogether, this implies the assertion. ∎
@@
Again, $Q_{\tau}^{\mathrm{av},T_{\max},h}$ is the same as the Besag-Clifford $p$-value $Q^{\mathrm{BC}}_{\gamma(h,T_{\max})}$ if we would only observe losses after step $\tau$; and if we would continue sampling after $\gamma(h,T_{\max})$, the p-value $Q_{\tau}^{\mathrm{av},T_{\max},h}$ would not change. This generalizes our results from Section 2.1, where we showed that the aggressive strategy recovers the Besag-Clifford method for $h=1$, to general $h$. Indeed, $Q_{t}^{\mathrm{av},\infty,1}=Q^{\mathrm{agg}}_{t}$ for all $t\in\mathbb{N}$.
@@
It is interesting how the anytime-valid versions of the permutation p-value $Q^{\mathrm{perm}}$ and the Besag-Clifford p-value $Q^{\mathrm{BC}}$ arise from our approach. Indeed, the betting function $E^{T}_{\alpha}$ in (17) is the most powerful bet we could make for fixed $\alpha$ and number of permutations $T$. Therefore, the resulting sequential betting strategy $B_{1}^{\alpha,T},\ldots,B_{T}^{\alpha,T}$ is optimal if $\alpha$ and $T$ are fixed. However, we claim that fixing the number of permutations is not the best way to perform permutation testing, since it is more reasonable to draw a larger number of permutations if the decision is tight and fewer permutations if the decision is unambiguous. This is provided by our mimicked log-optimal strategies.